OpenStudy (anonymous):

use the distributive property to simplify (-5-c)(-1)

OpenStudy (anonymous):

distributive property A(B+C)=AB+BC so (-5-c)*(-1)=(-5*-1)+(-c*-1)=5+c[as product of 2 -ve terms is +ve] :)